# Tidally Delayed Spin-Down of Very Low Mass Stars

### v2

@@ 4.3 Challenges for gyrochronology
Gyrochronology can in principle determine isolated low-mass MS stellar ages with reasonable precision (Barnes, 2003). But our study shows how the one-to-one mapping on which this method depends is broken with companions. Moreover, although we have explored scenarios with just one companion, the occurrence rate of planets around M dwarfs is at least 3.0 planets per star, potentially exacerbating the influence of tides on spin (Tuomi et al., 2019) and age-rotation degeneracy. Recalibrating gyrochronology to account for tides would be desirable to extend its utility beyond isolated stars. This is challenging but maybe possible because, for a large range of companion BD masses over a large age range, the period evolves little (see Fig. 2 ). Also, the curves for heavier companions are close together, and less sensitive to the influence of a specific companion mass.
@@ 5 Conclusions and Outlook
We have studied the combined influence of tides and magnetic spin-down of M-Dwarfs for different companion BD masses, initial orbital separations, stellar rotations and magnetic field strengths. We find that plausible M-Dwarf+BD binary interactions offer a possible explanation for the observed period distribution gap. Fig. 1, Fig. 2 and Fig. 3 reveal this gap differently. In Fig. 1, the spin-evolution tracks exhibit a period gap, followed by sharp spin-up spikes during tidal shredding, followed by steeply rising curves that result from the rapid M-dwarf spin-down after companion engulfment. Fig. 2 shows that the spin evolution tracks are separated by a period gap over their full age for most of the companion masses. In Fig. 3 the M-dwarfs initially spin down, but the companions move inward fast enough to rapidly spin-up the star, consistent with a period gap (E.g. Newton et al. (2016)). After companions are engulfed, the stars rapidly spin down by magnetic braking. Since the spin-down transition is fast, the fraction of stars therein is small and the overall period distribution becomes approximately bimodal with fast and slow rotator populations.
@@
For a $0.6M_{\odot}$ partially convective star however, Fig. 5 shows that a uniform distribution of companion masses produces more evenly spaced curves than the more pronounced slow and fast rotator population branches for fully convective M dwarf cases of Figs. 1, 2 and 3. This is independent of any magnetic topology changes or changes in core to envelope magnetic coupling, which have elsewhere been proposed to influence spin evolution (Garraffo et al., 2018; David et al., 2022; Cao et al., 2023). In our framework, the presence of a pronounced gap between slow and fast rotators for partially convective stars compared to fully convective M-dwarfs would reveal differences in companion mass and orbital separation distributions between the two stellar population classes, and may constrain the statistics of stellar-planetary system formation outcomes.
@@
Observations show that lower mass M dwarfs have a companion to primary mass ratio $m/M_{\star}$, that increases exponentially with decreasing mass for M-dwarfs of masses $M_{\star}\lesssim 0.25M_{\odot}$. In addition, massive M-dwarfs of $M_{\star}\gtrsim 0.35M_{\odot}$ lack companions at very close separations (Winters et al., 2019). These trends would imply statistically larger tidal torques and a larger gap between slow and fast rotators for fully convective M-dwarfs (see equation (10)), and relative to their partially convective counterparts.
@@
Knowing either companion mass or orbital separation from observations for a specific system can eliminate some degeneracy between the two in our model. The predicted spin evolution is more sensitive to orbital separation than to companion mass and so observations that pin down the orbital separation distributions will help to constrain the predicted fraction of M-dwarfs whose spin is tidally influenced. For example, a uniform distribution of companion separations could imply that in $\sim 80\%$ of systems the companion would not influence spin substantially (Fig. 4) whilst a higher fraction of companions at smaller separations would exacerbate the relative population of old M-dwarfs with anomalously short periods. Complementarily, the observed gap for old M-dwarfs can be used to constrain BD or planet mass-orbital separation distributions around fully convective M dwarfs. If tides cause the gap, this would signature the presence of BD companions and prioritize particular systems for companion searches in follow-up observations.
We have included only equilibrium tides in this work. Dynamical tides can enhance the tidal influence and offset the spindown of even more magnetically active M-dwarfs than we have considered. This would be particularly important for modeling specific M dwarf-BD system,s were data to become available.
@@
Finally, although observations show no sharp weakening of the rotation-activity at the fully convective M dwarf boundary (around M3 spectral type) they show a sharp drop in activity for stars later than spectral type M8 (Mohanty and Basri, 2003; Mohanty et al., 2002). (Mohanty et al., 2002) suggested that magnetic braking is less effective in stars later than M8 because a low ionization fraction leads to a high enough magnetic diffusivity such that the magnetic field decouples from the plasma near the surface. This implies a correlation between low activity and short spin periods in these M-dwarfs that is also consistent with our calculations, as tides do little to exacerbate spin down for these cases. We would then predict no significant population of long period low mass M-dwarfs later than $M8$, unless magnetic braking can survive in stars with these low ionization layers.

### v3

@@ 4.3 Challenges for gyrochronology
Gyrochronology can in principle determine isolated low-mass MS stellar ages with reasonable precision (barnes2003rotational). But our study shows how the one-to-one mapping on which this method depends is broken with companions. Moreover, although we have explored scenarios with just one companion, the occurrence rate of planets around M dwarfs is at least 3.0 planets per star, potentially exacerbating the influence of tides on spin (Tuomi2019) and age-rotation degeneracy. Recalibrating gyrochronology to account for tides would be desirable to extend its utility beyond isolated stars. This is challenging but maybe possible because, for a large range of companion BD masses over a large age range, the period evolves little (see Fig. 2 ). Also, the curves for heavier companions are close together, and less sensitive to the influence of a specific companion mass.
@@ 5 Conclusions and Outlook
We have studied the combined influence of tides and magnetic spin-down of M-Dwarfs for different companion BD masses, initial orbital separations, stellar rotations and magnetic field strengths. We find that plausible M-Dwarf+BD binary interactions offer a possible explanation for the observed period distribution gap. Fig. 1, Fig. 2 and Fig. 3 reveal this gap differently. In Fig. 1, the spin-evolution tracks exhibit a period gap, followed by sharp spin-up spikes during tidal shredding, followed by steeply rising curves that result from the rapid M-dwarf spin-down after companion engulfment. Fig. 2 shows that the spin evolution tracks are separated by a period gap over their full age for most of the companion masses. In Fig. 3 the M-dwarfs initially spin down, but the companions move inward fast enough to rapidly spin-up the star, consistent with a period gap (E.g. Newton2016). After companions are engulfed, the stars rapidly spin down by magnetic braking. Since the spin-down transition is fast, the fraction of stars therein is small and the overall period distribution becomes approximately bimodal with fast and slow rotator populations.
@@
For a $0.6M_{\odot}$ partially convective star however, Fig. 5 shows that a uniform distribution of companion masses produces more evenly spaced curves than the more pronounced slow and fast rotator population branches for fully convective M dwarf cases of Figs. 1, 2 and 3. This is independent of any magnetic topology changes or changes in core to envelope magnetic coupling, which have elsewhere been proposed to influence spin evolution (Garraffo2018; David2022; Cao2023). In our framework, the presence of a pronounced gap between slow and fast rotators for partially convective stars compared to fully convective M-dwarfs would reveal differences in companion mass and orbital separation distributions between the two stellar population classes, and may constrain the statistics of stellar-planetary system formation outcomes.
@@
Observations show that lower mass M dwarfs have a companion to primary mass ratio $m/M_{\star}$, that increases exponentially with decreasing mass for M-dwarfs of masses $M_{\star}\lesssim 0.25M_{\odot}$. In addition, massive M-dwarfs of $M_{\star}\gtrsim 0.35M_{\odot}$ lack companions at very close separations (Winters+2019). These trends would imply statistically larger tidal torques and a larger gap between slow and fast rotators for fully convective M-dwarfs (see equation (12)), and relative to their partially convective counterparts.
@@
Knowing either companion mass or orbital separation from observations for a specific system can eliminate some degeneracy between the two in our model. The predicted spin evolution is more sensitive to orbital separation than to companion mass and so observations that pin down the orbital separation distributions will help to constrain the predicted fraction of M-dwarfs whose spin is tidally influenced. For example, a uniform distribution of companion separations could imply that in $\sim 80\%$ of systems the companion would not influence spin substantially (Fig. 4) whilst a higher fraction of companions at smaller separations would exacerbate the relative population of old M-dwarfs with anomalously short periods. Complementarily, the observed gap for old M-dwarfs can be used to constrain BD or planet mass-orbital separation distributions around fully convective M dwarfs. If tides cause the gap, this would signature the presence of BD companions and prioritize particular systems for companion searches in follow-up observations. We have included only equilibrium tides in this work. Dynamical tides can enhance the tidal influence and offset the spindown of even more magnetically active M-dwarfs than we have considered. This would be particularly important for modeling specific M dwarf-BD systems were data to become available.
@@
Finally, although observations show no sharp weakening of the rotation-activity at the fully convective M dwarf boundary (around M3 spectral type) they show a sharp drop in activity for stars later than spectral type M8 (Mohanty+Basri2003Apj; Mohanty+2002ApJ). (Mohanty+2002ApJ) suggested that magnetic braking is less effective in stars later than M8 because a low ionization fraction leads to a high enough magnetic diffusivity such that the magnetic field decouples from the plasma near the surface. This implies a correlation between low activity and short spin periods in these M-dwarfs that is also consistent with our calculations, as tides do little to exacerbate spin down for these cases. We would then predict no significant population of long period low mass M-dwarfs later than $M8$, unless magnetic braking can survive in stars with these low ionization layers.
@@ Acknowledgments
Figure 6: Evolution of toroidal and radial magnetic fields for solving our dynamical equations in the top and bottom panels in red and blue respectively for $\epsilon=0.01$. For the initial stellar values, we used a mass of $M=0.2M_{\star}$, a rotation period $P_{\star}=0.1$ days, and a magnetic field $B_{\rm p,\star i}=50\rm G$. At each point, on these curves, the tangent lines would represent a power low scalings of the form $b_{\phi}(t)\propto t^{\xi_{1}(t)}$ and $b_{r}(t)\propto t^{\xi_{2}(t)}$. The black lines represent approximations of single best fit line that approximates $\xi_{1}(t),\xi_{2}(t)$ from $t\approx 1.5-10$ Gyr. From these line scalings shown, we solve for $b_{\phi}(t)$ as a function of $b_{r}(t)$ and further account for $|B_{r}(t)|\propto b_{r}(t)^{3}$ to infer $\eta=\xi_{1}/(3\xi_{2})$ for relation between the toroidal field and predicted observed radial magnetic field $b_{\phi}(t)\propto C|B_{r}(t)|^{\eta}$. In this case $\eta=0.96.$. This rough scaling is reasonably consistent with observed ranges (See2015).